OpenStudy (anonymous):

Whats The Decimal Of 3/6

OpenStudy (anonymous):

3/6 = 1/2, 1/2 =?

OpenStudy (anonymous):

3/6 is reduced to 1/2 and 1/2 equals to 0.5